Project

General

Profile

05282014 Minutes

Attending: Barnali, Nitin, Gavin, Michael and Mat

Nitin: Evan sent categories of the plots.
- Micheal name should not be the contact for the webpages anymore.

FD:

- Lots of computing center over the weekend
- 9 am funny jump in FEB hit rate spectrum and average number of hits per spill. Many FEBs change rates.
- 7:30 drop in number of active FEBs (recovered by sync).
- P1 4-11 as of 5/27
Percent of Empty spill: 0.045, once every couple hour one event is empty
Negative time jump reported by message analyzer. (Caused by hot channels?)

Near Det -
Working on the new timing plot. Might show up today or tomorrow.

Last NUMI reco through the end of April.